A 5.0 kg bucket of water is raised from a well by a rope. If the upward accelaration of the bucket is 3.0m/s2 find the force exerted by the rope on the bucket


1
Expert's answer
2021-11-14T17:17:43-0500

Given:

"m=5.0 \:\rm kg"

"a=3.0 \:\rm m/s^2"


The Newton's second law says

"ma=F-mg"

Hence

"F=m(a+g)=5.0*(3.0+9.8)=64\:\rm N"